the united states has minted one dollar silver coins called the american eagle silver bullion coins since 1986. Each coin has a diameter of 40.6 millimeters and is 2.98 millimeters thick. the density of silver is 10.5 grams per cubic centimeter. what is the mass of an american eagle silver bullion coin to the nearest gram?
Answer by Boreal(15235)

Volume is pi*r^2*thickness=pi* 412.09*2.98=3857.96 mm^3
1000 mm^3=1 cm^3
Therefore, 3.858 cm^3
multiply that by 10.5 g/cm^3 to get grams
41 grams.